### v2

# PyPLUTO: a data analysis Python package for the PLUTO code

[…]
PyPLUTO: a data analysis Python package for the PLUTO code
Giancarlo Mattia  Max-Planck-Institut für Astronomie, Königstuhl 17, Heidelberg, 69117, Germany Daniele Crocco  Dipartimento di Fisica, Università di Torino, Via P. Giuria 1, Torino, 10125, Italy David Melon Fuksman  Max-Planck-Institut für Astronomie, Königstuhl 17, Heidelberg, 69117, Germany Matteo Bugli  Dipartimento di Fisica, Università di Torino, Via P. Giuria 1, Torino, 10125, Italy Institut d’Astrophysique de Paris, UMR 7095, CNRS & Sorbonne Université, 98 bis boulevard Arago, 75014 Paris, France Université Paris-Saclay, Université Paris Cité, CEA, CNRS, AIM, Gif-sur-Yvette, 91191, France INFN, Sezione di Torino, Via P. Giuria 1, Torino, 10125, Italy Vittoria Berta  Dipartimento di Fisica, Università di Torino, Via P. Giuria 1, Torino, 10125, Italy INFN, Sezione di Torino, Via P. Giuria 1, Torino, 10125, Italy Eleonora Puzzoni  Observatoire de la Côte d’Azur, Laboratoire Lagrange, Bd de l’Observatoire, CS 34229, 06304 Nice cedex 4, France Andrea Mignone  Dipartimento di Fisica, Università di Torino, Via P. Giuria 1, Torino, 10125, Italy INFN, Sezione di Torino, Via P. Giuria 1, Torino, 10125, Italy Bhargav Vaidya  Department of Astronomy, Astrophysics and Space Engineering, Indian Institute of Technology, Khandwa Road, Simrol, Indore, 453552, India
(September 17, 2025)
Keywords: Astronomy, Python, MagnetoHydroDynamics, Computational Astrophysics, Data Visualization
Summary
In recent years, numerical simulations have become indispensable for addressing complex astrophysical problems. The so-called magnetohydrodynamics (MHD) framework represents a key tool for investigating the dynamical evolution of astrophysical plasmas. This formalism consists of a set of partial differential equations (Chiuderi & Velli, 2015) that enforce the conservation of mass, momentum, and energy, along with Maxwell’s equations for the evolution of the electromagnetic fields. Due to the high nonlinearity of the MHD equations (regardless of their specifications, e.g., classical/relativistic or ideal/resistive), a general analytical solution is not possible, making numerical approaches crucial. Numerical simulations usually produce large sets of data files, and their scientific analysis relies on dedicated software tools designed for data visualization (Ahrens et al., 2005; Childs et al., 2012). However, to encompass all code output features, specialized tools focusing on the numerical code may represent a more versatile and integrated solution.
Here, we present PyPLUTO, a Python package tailored for efficient loading, manipulation, and visualization of outputs produced with the PLUTO [^1] https://plutocode.ph.unito.it code (Mignone et al., 2007; Mignone, Zanni et al., 2012). PyPLUTO uses memory mapping to optimize data loading and provides general data manipulation and visualization routines. PyPLUTO also supports the particle modules of the PLUTO code, enabling users to load and visualize particles, such as cosmic rays (Mignone et al., 2018), Lagrangian (Vaidya et al., 2018), or dust particles (Mignone et al., 2019), from hybrid simulations. A dedicated graphical user interface (GUI, shown in Fig. 1) simplifies the generation of single-subplot figures, making PyPLUTO a powerful yet user-friendly toolkit for astrophysical data analysis.
Figure 1: Interactive visualization of shock tube test results (i.e., density, pressure, and velocity profiles) with the GUI.
Statement of need
The PLUTO code (Mignone et al., 2007) is a widely-used, freely-distributed computational fluid dynamics code designed to solve the classical and (special) relativistic MHD equations in different geometries and spatial dimensions. The original code is written in C (while the upcoming GPU version provides a complete C++ rewrite [^2] https://plutocode.ph.unito.it/pluto-gpu.html) and it contains several numerical methods adaptable to different contexts. Data post-processing is a crucial step in analyzing the results of any numerical simulation. Other packages addressing related needs (e.g plutoplot [^3] https://github.com/Simske/plutoplot), provide valuable functionality for working with PLUTO data, including loading and visualization. However, they may not support all data formats or offer integration for data manipulation and advanced plotting tasks. In this work, we present PyPLUTO, a Python package designed to efficiently load, manipulate, and visualize the output from the PLUTO code.
The initial version of PyPLUTO (written by Bhargav Vaidya) is available here [^4] https://github.com/coolastro/pyPLUTO. PyPLUTO has since been completely rewritten and is now maintained at a new repository [^5] https://github.com/GiMattia/PyPLUTO. The package retains its core strengths while offering user-friendly methods for generating publication-quality plots with high customization. In addition to its enhanced flexibility, PyPLUTO offers strong computational efficiency, enabling the rapid handling of large datasets typical of state-of-the-art numerical simulations. Through this balance between customization, performance, and ease of use, PyPLUTO represents a key tool to effectively communicate scientific results while minimizing the effort required for post-processing.
Main features
PyPLUTO is a package written for Python 3.10 or later with the additions of external packages such as NumPy (Harris et al., 2020), Matplotlib (Hunter et al., 2007), SciPy (Virtanen et al., 2020), pandas (McKinney, 2010), h5py (Collette, 2013), and PyQT6666https://www.riverbankcomputing.com/software/pyqt/intro. The package, which can be installed through pip, primarily consists of three main classes:
  • The Load class loads and manipulates the PLUTO output files containing fluid-related quantities.
  • The LoadPart class loads and manipulates the PLUTO output files containing particle-related quantities.
  • The Image class produces and handles the graphical windows and the plotting procedures.
Additionally, a separate PyPLUTOApp class launches a GUI able to load and plot 1D and 2D data in a single set of axes. PyPLUTO has been implemented to be supported by Windows, macOS, and Linux, through both standard scripts and more interactive tools (e.g., IPython or Jupyter). The style guidelines follow the PEP8777https://peps.python.org/pep-0008/ conventions for Python codes, enforced through the Black package Langa et al., 2020, and focus on clarity and code readability. Finally, by leveraging the capabilities of the Sphinx package [^8] http://sphinx-doc.org/sphinx.pdf, PyPLUTO features extensive docstrings, providing a helpful reference for both users and developers.
Benchmark examples
PyPLUTO provides a set of benchmarks that are immediately accessible after the package is installed. These consist of test problems that can be applied to relevant astrophysical applications and showcase the full range of PyPLUTO’s features. Here we report two examples demonstrating the package’s capabilities.
Disk-Planet Interaction
This test simulates the interaction of a planet embedded in a disk (Mignone, Flock et al., 2012) and represents an ideal scenario for understanding the formation and evolution of planetary systems. In particular, the formation of spiral density waves and disk gaps represents some key observational signatures of planet formation and planet-disk interaction (Melon Fuksman et al., 2021, Muley et al., 2024). In the left panel of Fig. 2, we show an adaptation of Figure 10 of Mignone, Flock et al., 2012, featuring two separate zoom-ins around the planet’s location.
  • The first zoom (upper-right subplot) shows an enlarged view of the density distribution using the same color map and logarithmic scale as the global plot.
  • The second zoom (lower-left subplot) highlights the changes in toroidal velocity due to the planet’s presence by employing a different color map (to enhance the sign change) and a linear color scale.
Figure 2: Left panel: example of inset zooms of the planet region of the disk-planet test problem. The main plot and the right zoom show the density on a logarithmic scale, while the left zoom highlights the toroidal velocity on a linear scale. Right panel: example of an X-point region with magnetic field lines overlaid (as contour lines of the vector potential, solid lines). The main plot shows the test-particle distribution, color-coded by velocity magnitudes, while the inset plot displays the particle energy spectrum at the beginning (in blue) and end (in red) of the simulation.
These zoomed views offer deeper insights into the physical processes at play and demonstrate the utility of PyPLUTO for analyzing complex astrophysical systems.
Particles Accelerated near an X-point
This test problem examines particle acceleration near an X-type magnetic reconnection region (Puzzoni et al., 2021). In the last decades, magnetic reconnection (Mattia et al., 2023, Bugli et al., 2024) has proven to be a key physical process to explain the population of non-thermal particles in solar flares, relativistic outflows, and neutron star magnetospheres. This sort of test provides valuable insights into particle acceleration mechanisms in high-energy astrophysical environments by enabling the investigation of particle trajectories and energy distribution near the X-point.
In the right panel of Fig. 2 we show an adaptation of the top panel of Figure 13-14 from Mignone et al., 2018. The main plot displays the distribution of test particles, color-coded by their velocity magnitudes, with magnetic field lines overlaid as solid and dashed lines. The inset panel shows the energy spectrum at the initial ($t=0$, in blue) and final ($t=100$, in red) time. In this scenario, the absence of a guide field ($\mathbf{E}\cdot\mathbf{B}=0$) results in a symmetric distribution along the y-axis from the combined effects of the gradient, curvature, and $\mathbf{E}\times\mathbf{B}$ drifts in the vicinity of the X-point, where the electric field is the strongest.
This plot provides a clear visual representation of particle motion and energy changes, demonstrating how PyPLUTO can be used to investigate complex processes such as particle acceleration in astrophysical sources.
Ongoing research using PyPLUTO
Research applicable with PyPLUTO includes the development of numerical algorithms (Mattia & Mignone, 2022, Berta et al., 2024, Melon Fuksman et al. 2025) and numerical simulations of astrophysical objects, such as jets (Mattia & Fendt, 2022, Mattia et al. 2023, Mattia et al. 2024, Costa et al. 2025, Sciaccaluga et al. 2025), star clusters (Härer et al. 2025) and protoplanetary disks (Melon Fuksman et al. 2024a, Melon Fuksman et al. 2024b), as well as physical processes, such as particle acceleration (Wang et al., 2024) and magnetic reconnection (Bugli et al., 2025).
Conclusion and Future Perspectives
The PyPLUTO package is designed as a powerful yet flexible tool to facilitate the data analysis and visualization of the output from PLUTO simulations, focusing on user friendliness while allowing the necessary customization to produce publication-quality figures. To overcome current limitations and further enhance the package’s capabilities, particular focus will be devoted to:
  • introducing specific routines for rendering 3D data to provide users with tools for visualizing volumetric data;
  • supporting interactive visualization and comparison of multiple simulation outputs, allowing the users to track temporal evolution directly with the GUI; and
  • expanding the graphical interface to support particle data, including dynamic visualization of particle distributions and trajectories.
Alongside these improvements, PyPLUTO development will focus on encompassing the latest features of the PLUTO code, such as new adaptive mesh refinement strategies and extensions to more general metric tensors. PyPLUTO is a public package that can be downloaded alongside the CPU and GPU versions of the PLUTO code [^9] https://gitlab.com/PLUTO-code/gPLUTO. Regular updates will be released with improvements and bug fixes. Additionally, a repository [^10] https://github.com/GiMattia/PyPLUTO containing the PyPLUTO development versions will be available for users who wish to exploit the code’s latest features in advance.
Acknowledgments
The authors thank the reviewers for improving this work with valuable comments and suggestions. G. Mattia thanks L. Del Zanna and M. Flock for the discussions on data visualization and the Data Science Department of the Max Planck Institute for Astronomy for helping with Python and Matplotlib.
The authors thank Simeon Doetsch for their insights on memory-mapping techniques and Deniss Stepanovs and Antoine Strugarek for contributing to previous PyPLUTO versions throughout the years. The authors thank Agnese Costa, Alberto Sciaccaluga, Alessio Suriano, Asmita Bhandare, Dhruv Muley, Dipanjan Mukherjee, Jieshuang Wang, Jacksen Narvaez, Lucia Haerer, Prakruti Sudarshan, Stefano Truzzi, and Stella Boula for testing the module while it was under development. M. Bugli acknowledges the support of the French Agence Nationale de la Recherche (ANR), under grant ANR-24-ERCS-0006 (project BlackJET). This project has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101064953 (GR-PLUTO), and from the European High Performance Computing Joint Undertaking (JU) and Belgium, Czech Republic, France, Germany, Greece, Italy, Norway, and Spain under grant agreement No 101093441 (SPACE).